OpenStudy (anonymous):

The dimensions of Eric’s pool are 9 feet by 7 feet. He wants to build a walkway surrounding the pool so that the square footage of the walkway and pool is 143 square feet. How wide is one side of the walkway?

OpenStudy (a_clan):

Area of walkway = 143 - 63 = 80 sq ft So, 2[(9+2x)*x] + 2[7*x] = 80

OpenStudy (a_clan):

It is not allowing me to attach the image file